OpenStudy (amyna):

Detrmine monotonicity: (0.9)^n How do i solve this?

zepdrix (zepdrix):

(0.9)^1 = 0.9 (0.9)^2 = 0.81 (0.9)^3 = 0.729 It's clearly monotonically decreasing as every term is smaller than the one before it. But how to show this -_- Hmm this calc 2? I don't remember the formulas lol

OpenStudy (amyna):

yes its calc 2, i know its decreasing, but i have to prove it. My teacher said there are 2 ways: 1) is to take the an+1 sequence thing 2) is to take the derivative

zepdrix (zepdrix):

If the sequence is decreasing, then the ratio\[\large\rm \frac{a_{n+1}}{a_n}\lt 1.\]Because the numerator should be `smaller` if it's decreasing, ya?

OpenStudy (amyna):

yes

zepdrix (zepdrix):

\[\large\rm \frac{a_{n+1}}{a_n}=\frac{0.9^{n+1}}{0.9^n}=0.9\lt1\qquad \forall n\in\mathbb N\]So maybe we can do just this?

OpenStudy (amyna):

yes this seems correct! thank you!

zepdrix (zepdrix):

sequences stuff is tricky! >.<

OpenStudy (amyna):

ya i know lol!